cos x + cos y =1/3, sin x + sin y =1/4 then sin(x+y)
a.25/7
b.24/25

Grade:11

1 Answers

Deepak Kumar Shringi
5 years ago


2 sin x+y/2 cos x-y/2 = 1/3
&
2 cos x+y/2 cos x-y/2 = ¼

Now divide
tan (x+y/2) = 4/3

Now

Sin(x+y) = 2 tan (x+y/2) /1 + tan² (x+y/2)

= (8/3) /(1 + 16/9)

= 24/25